40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 23 20:32:43 2015, Koji, Update, LSC, Frequency source fixed, IMC LO level adjusted 
    Reply  Mon Nov 23 21:48:49 2015, Koji, Update, LSC, Frequency source fixed, IMC LO level adjusted IMG_2136.JPGIMG_2129.JPGIMG_2133.JPG
       Reply  Thu Nov 26 19:39:27 2015, Koji, Update, LSC, Current state of the frequency source, and possible improvement RF_Frequency_Source.pdf
          Reply  Sun Nov 29 05:23:57 2015, rana, Update, LSC, Current state of the frequency source, and possible improvement 
             Reply  Sun Nov 29 12:32:26 2015, Koji, Update, LSC, Current state of the frequency source, and possible improvement 
Message ID: 11822     Entry time: Sun Nov 29 12:32:26 2015     In reply to: 11821
Author: Koji 
Type: Update 
Category: LSC 
Subject: Current state of the frequency source, and possible improvement 

I'm not claiming we need to modify the frequency source immediately as we are not limited by the oscillator amplitude or phase noise.
I just wanted to note something in mind before it goes away quickly.

Alberto's T1000461 tells us that the oscillator and phase noise are degraded by factor of ~3 and ~5 due to the RF chanin.
My diagram is possible removal of up-down situation of the chain.

Maybe more direct improvement would be:

- Removal of two amplifiers out of four. The heat condition of the box is touch thought it is not critical.

- The modification will allow us to have a spare 11MHz channel at 1X2 rack that would be useful for 3f modulation.

ELOG V3.1.3-